A salt has a formula Na2MO3. The mass of 0.225 mol of this compound is 23.8 g. What is the identity of M?

A. Be

B. B

C. C

D. N

E. Cl
Answer #1

Say molar mass of M = y g/mol. Molar mass of Na is 23 g/mol and that of O is 16 g/mol.

So, molar mass of the salt = (2x23 + y + 3x16) g/mol = (94+y) g/mol.

We know, molar mass = mass of 1 mole of substance

Again, 0.225 mol of the salt has a mass 23.8 g

Thus, 1 mol of it has a mass (23.8 / 0.225) g = 105.78 g = 106 g (approximately)

So,

94 + y = 106

Or, y = 12

Element having molar mass 12 g/mol is C.

So, M = C (C is the correct option)